### v1

# Convergence analysis of Levenberg-Marquardt method with Singular Scaling for non-zero residue nonlinear least-squares problems [^1] This work is supported by FAPESC (Fundação de Amparo à Pesquisa e Inovação do Estado de Santa Catarina) [grant number 2023TR000360]

Abstract
This paper addresses the convergence analysis of a variant of the Levenberg-Marquardt method (LMM) designed for nonlinear least-squares problems with non-zero residue. This variant, called LMM with Singular Scaling (LMMSS), allows the LMM scaling matrix to be singular, encompassing a broader class of regularizers, which has proven useful in certain applications. In order to establish local convergence results, a careful choice of the LMM parameter is made based on the gradient linearization error, dictated by the nonlinearity and size of the residual. Under completeness and local error bound assumptions we prove that the distance from an iterate to the set of stationary points goes to zero superlinearly and that the iterative sequence converges. Furthermore, we also study a globalized version of the method obtained using line-search and prove that any limit point of the generated sequence is stationary. Some examples are provided to illustrate our theoretical results.
Keywords. Levenberg-Marquardt, Singular Scaling Matrix, Convergence analysis, Non-zero residue
1 Introduction
In this paper, we investigate global and local convergence properties of the Levenberg–Marquardt method applied to a nonlinear least-squares (NLS) problem:
$\min_{x\in\mathbb{R}^{n}}\quad\frac{1}{2}\|F(x)\|^{2}:=\phi(x),$ (1)
where $F:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ is twice continuously differentiable. In particular, we focus on the overdetermined problem, where $m\geq n$.
Unlike previous research, which predominantly focused on the zero residue case or in cases where the rank of the Jacobian is complete [7, 6, 10, 11, 13], here we consider NLS problems where the set of zeros of $F$ is empty and its Jacobian is allowed to be rank-deficient.
We say that problem (1) is a non-zero-residue NLS problem when the set of zeros of $F$ is empty. Our interest is to identify the stationary points of $\phi$, i.e., the set $X^{*}=\{x\in\mathbb{R}^{n}\mid J(x)^{T}F(x)=0\}$, where $J(x)\in\mathbb{R}^{m\times n}$ is the Jacobian of $F$ at $x$, under the assumption that $X^{*}\neq\emptyset$. Throughout the text we shall use $x^{*}$ to denote an arbitrary element of $X^{*}$.
Although works such as [2, 3] explore problems with non-zero residuals and present results of local convergence of the LM method, it should be noted that the scale matrix used in these works is the identity matrix. In this work, in addition to assuming non-zero residue, we consider the following iteration:
$(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)d_{k}$ $=-J_{k}^{T}F_{k}$ (2)
$x_{k+1}$ $=x_{k}+\alpha_{k}d_{k},\quad\forall k\geq 0,$ (3)
where $F_{k}:=F(x_{k})$, $J_{k}:=J(x_{k})$ is the Jacobian of $F$ at $x_{k}$, $\lambda_{k}>0$, $\alpha_{k}$ is the step size, and $L^{T}L$ referred to as scaling matrix, is allowed to be singular. We refer to the iteration (2)–(3) as the Levenberg-Marquardt method with Singular Scaling (LMMSS).
This type of iteration is motivated by applications in inverse problems, as recently seen in [5], where a parameter recovery problem using LMMSS is addressed, including a local convergence analysis for the zero residue case. On the other hand, convergence properties of the classic LMM in the non-zero residue scenario have received recent attention in the literature [2, 3]. These are two aspects in which our contribution differs from the existing literature.
Our aim is to analyze the local convergence of LMM with singular scaling matrices in the non-zero residue scenario. Furthermore, we shall provide a globalization for the LMMSS iteration where the step-size $\alpha_{k}$ is determined by line-search and prove that any limit point of the generated sequence is stationary for the least-squares function.
This paper is organized as follows. In Section 2, we start with some assumptions that will be considered throughout the manuscript. In Section 3, we present technical results necessary for the upcoming sections. We show that LMMSS iteration is well-defined and provide the local convergence results in Section 4, which are categorized into two cases based on the behavior of the rank of the Jacobian around a stationary point (constant and diminishing rank). Finally, Section 5 proves that, for LMMSS with an Armijo line-search scheme, every limit point of the generated sequence is a stationary point of the sum-of-squares function, thereby establishing the global convergence of this method. A few examples are given in Section 6 to illustrate our theoretical results. Final considerations are given in Section 7.
2 Assumptions
Recall that throughout this manuscript, we assume that the set of stationary points of $\phi$ is non-empty: $X^{*}\neq\emptyset$.
Assumption 1 below guarantees that $d_{k}$ solution of (2) is unique.
Assumption 1.
The matrix $L\in\mathbb{R}^{p\times n}$ has $\text{rank}(L)=p\leq n\leq m$ and there exists $\Omega\subset\mathbb{R}^{n}$ and $\gamma>0$ such that, for every $x\in\Omega$
$\left\lVert J(x)v\right\rVert^{2}+\left\lVert Lv\right\rVert^{2}\geq\gamma\left\lVert v\right\rVert^{2},\quad\forall v\in\mathbb{R}^{n}.$ (4)
Such assumption comes from the inverse problems literature [12, 8] where it is sometimes called completeness condition. It was a key assumption in [5] to study the local convergence of LMMSS in the zero residue setting. We remark that such assumption is equivalent to $\mathcal{N}(J(x))\cap\mathcal{N}(L)=\{{\bf 0}\},\forall x\in\Omega.$
Next assumption is quite standard in the Gauss–Newton and Levenberg–Marquardt literature and asks the Jacobian to be locally Lipschitz.
Assumption 2.
For any $x^{*}\in X^{*}$, there exists a constant $\delta\in(0,1)$ and $L_{0}>0$ such that
$\|J(x)-J(y)\|\leq L_{0}\|x-y\|,$ (5)
for all $x,y\in B(x^{*},\delta)$.
Assumption 2 implies that:
$\|J(y)(x-y)-(F(x)-F(y))\|\leq L_{1}\|x-y\|^{2},\quad\forall x,y\in B(x^{*},\delta)$ (6)
where $L_{1}=L_{0}/2$, that is, the error in the linear approximation of $F(x)$ around $y$ is $O(\|x-y\|^{2})$, for $x$ and $y$ in a neighborhood of $x^{*}$.
Due to the compactness of the ball $B(x^{*},\delta)$, there exist positive constants $L_{2}$ and $\beta$ such that $\|J(x)\|\leq L_{2}$ and $\|F(x)\|\leq\beta$ for all $x\in B(x^{*},\delta)$. Therefore, since $\|J(x)\|$ is bounded in $B(x^{*},\delta)$, by applying the mean value inequality, we can infer that $\|F(x)-F(y)\|\leq L_{2}\|x-y\|$, for all $x,y\in B(x^{*},\delta)$. Additionally, the gradient $\nabla\phi(x)=J(x)^{T}F(x)$ is Lipschitz in $B(x^{*},\delta)$:
$\left\lVert J(x)^{T}F(x)-J(y)^{T}F(y)\right\rVert\leq L_{3}\|x-y\|,\quad\forall x,y\in B(x^{*},\delta),$ (7)
where $L_{3}=L_{2}^{2}+\beta L_{0}.$
Moreover, notice that for $z\in X^{*}\cap B(x^{*},r)$ and $x,y\in B(x^{*},r)$, we have
$\|(J(x)-J(y))^{T}F(y)\|$ $=\|(J(x)-J(z)+J(z)-J(y))^{T}F(y)\|$ (8)
$\leq\|(J(x)-J(z))^{T}F(y)\|+\|(J(z)-J(y))^{T}F(y)\|$
$\leq L_{0}L_{2}\|x-z\|\|y-z\|+\|J(x)^{T}F(z)\|$
$+L_{0}L_{2}\|y-z\|^{2}+\|J(y)^{T}F(z)\|.$
Lemma 1.
[2, Lemma 2.1] If Assumption 2 is satisfied, then there exists a constant $L_{4}>0$ such that
$\left\lVert\nabla\phi(y)-\nabla\phi(x)-J(x)^{T}J(x)(y-x)\right\rVert\leq L_{4}\left\lVert x-y\right\rVert^{2}+\left\lVert(J(x)-J(y))^{T}F(y)\right\rVert,$ (9)
for all $x,y\in B(x^{*},\delta)$.
Next, we present the local error bound condition assumed in this work. We remark that it is an error bound condition with respect to the set of stationary points of $\phi$ rather than the zeroes of $F$ (which is empty in the case of non-zero residue).
Assumption 3 (Error bound).
For any $x^{*}\in X^{*}$, $\left\lVert J(x)^{T}F(x)\right\rVert$ provides an error bound in $B(x^{*},\delta)$, i.e., there exists $\omega\in(0,\infty)$ such that
$\omega\textup{dist}(x,X^{*})\leq\left\lVert J(x)^{T}F(x)\right\rVert,\quad\forall x\in B(x^{*},\delta),$ (10)
where $\textup{dist}(x,X^{*})=\inf_{z\in X^{*}}{\left\lVert x-z\right\rVert}$.
Throughout the text, given $x$, we shall denote by $\bar{x}$ an element of $X^{*}$ such that $\|x-\bar{x}\|=\text{dist}(x,X^{*})$. Also, let $x^{*}\in X^{*}$ and $J_{*}=J(x^{*})$.
From Assumptions 3 and (7), we obtain
$\omega\text{dist}(x_{k},X^{*})\leq\left\lVert J_{k}^{T}F_{k}\right\rVert\leq L_{3}\text{dist}(x_{k},X^{*}).$ (11)
The remaining assumption focuses, as outlined in (8), on the terms
$\left\lVert J(x)^{T}F(z)\right\rVert\text{ and }\left\lVert J(y)^{T}F(z)\right\rVert.$
These terms play a crucial role in controlling the error of the “incomplete linearization” of the gradient, as expressed in (9[2].
Assumption 4.
For every $x\in B(x^{*},\delta)$ and every $z\in X^{*}\cap B(x^{*},\delta)$, the following inequality holds:
$\|(J(x)-J(z))^{T}F(z)\|\leq C\|x-z\|^{1+r},$ (12)
with $r\in(0,1]$ and $C\geq 0$.
3 Auxiliary Results
In this section, we organize other fundamental results essential for the upcoming analysis.
We start by mentioning that (2) always has a solution. In fact, (2) corresponds to the normal equation $B^{T}Bd=B^{T}c$ of the augmented system
$Bd:=\left[\begin{array}[]{c}J_{k}\\ \sqrt{\lambda_{k}}L\end{array}\right]d=-\left[\begin{array}[]{c}F_{k}\\ 0\end{array}\right]=:c,$
and the normal equation always has a solution because $B^{T}c$ lies in the range of $B^{T}B$. But it is Assumption 1 that ensures the solution of (2) is unique. Indeed, it is easy to see that condition (4) at $x_{k}$ implies that $J_{k}^{T}J_{k}+\lambda_{k}L^{T}L$ is positive definite (and thus nonsingular).
Next result will be important when the rank of $J(x)^{T}J(x)$ is constant in a neighborhood of $x^{*}$.
Lemma 2.
[2, Corollary 3.1] Given $\kappa>1$, if
$\text{rank}(J(x)^{T}J(x))=\text{rank}(J_{*}^{T}J_{*})=q\geq 1,$
and
$\left\lVert J(x)^{T}J(x)-J_{*}^{T}J_{*}\right\rVert\leq\Big{(}1-\frac{1}{\kappa}\Big{)}\frac{1}{\left\lVert(J_{*}^{T}J_{*})^{+}\right\rVert},$ (13)
where $(J_{*}^{T}J_{*})^{+}$ denote the Moore-Penrose pseudo-inverse, then
$\left\lVert(J(x)^{T}J(x))^{+}\right\rVert\leq\kappa\left\lVert(J_{*}^{T}J_{*})^{+}\right\rVert.$
Now, we will show some results on the Generalized Singular Value Decomposition (GSVD), an important tool in theoretical analysis of the LMMSS direction.
Theorem 1.
(GSVD) [9, p. 22] Consider the pair $(A,L)$, where $A\in\mathbb{R}^{m\times n}$, $L\in\mathbb{R}^{p\times n}$, $m\geq n\geq p$, $rank(L)=p$, and $\mathcal{N}(A)\cap\mathcal{N}(L)=\{0\}$. Then, there exist matrices $U\in\mathbb{R}^{m\times n}$ and $V\in\mathbb{R}^{p\times p}$ with orthonormal columns and a nonsingular matrix $X\in\mathbb{R}^{n\times n}$ such that
$A=U\left[\begin{array}[]{cc}\Sigma&0\\ 0&I_{n-p}\end{array}\right]X^{-1}\quad\text{and}\quad L=V\left[\begin{array}[]{cc}M&0\end{array}\right]X^{-1},$ (14)
with $\Sigma$ and $M$ being the following diagonal matrices:
$\Sigma=\text{diag}(\sigma_{1},\dots,\sigma_{p})\in\mathbb{R}^{p\times p}\quad\text{and}\quad M=\text{diag}(\mu_{1},\dots,\mu_{p})\in\mathbb{R}^{p\times p}.$ (15)
Moreover, the elements of $\Sigma$ and $M$ are non-negative, ordered as follows:
$0\leq\sigma_{1}\leq\dots\leq\sigma_{p}\leq 1\quad\text{and}\quad 1\geq\mu_{1}\geq\dots\geq\mu_{p}>0,$ (16)
and normalized by the relation $\sigma_{i}^{2}+\mu_{i}^{2}=1$, for $i=1,\dots,p$. We call the generalized singular value of the pair $(A,L)$ the ratio
$\gamma_{i}=\frac{\sigma_{i}}{\mu_{i}},\quad i=1,\dots,p.$
Remark 1.
Since $X^{-1}$ is nonsingular and $U$ has orthonormal columns we have
$\text{rank}(A)=\text{rank}(U\left[\begin{array}[]{cc}\Sigma&0\\ 0&I_{n-p}\end{array}\right]X^{-1})=\text{rank}(\left[\begin{array}[]{cc}\Sigma&0\\ 0&I_{n-p}\end{array}\right])\geq n-p.$
By considering the GSVD of the pair $(J_{k},L)$ we can provide a useful characterization of the direction $d_{k}$ in LMMSS. In fact, given the GSVD
$J_{k}=U_{k}\left[\begin{array}[]{cc}\Sigma_{k}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{-1}\quad\text{and}\quad L=V_{k}\left[\begin{array}[]{cc}M_{k}&0\end{array}\right]X_{k}^{-1},$
where
$(\Sigma_{k})_{ii}:=\sigma_{i,k}\quad\text{and}\quad(M_{k})_{ii}:=\mu_{i,k},\quad i=1,\dots,p;$ (17)
it follows that
$J_{k}^{T}J_{k}+\lambda_{k}L^{T}L=X_{k}^{-T}\left[\begin{array}[]{cc}\Sigma_{k}^{2}+\lambda_{k}M_{k}^{2}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{-1}.$ (18)
Then, $d_{k}$ from (2) can be expressed as:
$d_{k}=-X_{k}\left[\begin{array}[]{cc}\Gamma_{k}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{T}{J_{k}}^{T}F_{k},$ (19)
with $\Gamma_{k}:=(\Sigma_{k}^{2}+\lambda_{k}M_{k}^{2})^{-1}.$
Next result from [5] gives an upper bound on $\|X_{k}\|$ depending on $\gamma$ from Assumption 1.
Lemma 3.
[5, Lemma 2.2] Consider $\{x_{k}\}$ the sequence generated by LMMSS along with the GSVD of the pair $(J_{k},L)$. Then, $\forall k\in\mathbb{N}$, $\|X_{k}\|\leq 1/\sqrt{\gamma}$.
Remark 2.
By the GSVD for the pair $(J_{k},L)$, it follows that
$J_{k}^{T}J_{k}+L^{T}L=X_{k}^{-T}X_{k}^{-1}.$
Therefore,
$\left\lVert X_{k}^{-1}\right\rVert^{2}\leq\left\lVert J_{k}^{T}J_{k}\right\rVert+\left\lVert L\right\rVert^{2}\leq L_{2}^{2}+\left\lVert L\right\rVert^{2}.$ (20)
Now, let us analyze an upper bound for $\left\lVert\Gamma_{k}\right\rVert$. From the GSVD for the pair $(J_{k},L)$, using the generalized singular values $\gamma_{i,k}=\sigma_{i,k}/\mu_{i,k}$, we have:
$\sigma_{i,k}^{2}=\frac{\gamma_{i,k}^{2}}{\gamma_{i,k}^{2}+1}\quad\text{and}\quad\mu_{i,k}^{2}=\frac{1}{\gamma_{i,k}^{2}+1}.$
Thus, we can express $\Gamma_{k}$ from (19) as:
$(\Gamma_{k})_{ii}=\frac{1}{\sigma_{i,k}^{2}+\lambda_{k}\mu_{i,k}^{2}}=\frac{\gamma_{i,k}^{2}+1}{\gamma_{i,k}^{2}+\lambda_{k}},\quad i=1,\dots,p.$
Therefore, we can bound $\|\Gamma_{k}\|$ in terms of the generalized singular values $\gamma_{i,k}$ and the parameter $\lambda_{k}$. To this end, we can use the following lemma.
Lemma 4.
For the function
$\psi(\gamma,\lambda)=\dfrac{\gamma^{2}+1}{\gamma^{2}+\lambda},\quad\gamma\geq 0,\ \lambda>0,$
the following properties hold
  • (a)
    For $\lambda\in(0,1)$, the function $\psi(\gamma,\lambda)$ has a unique maximum value attained at
    $\gamma_{\max}=0\quad\text{ where }\quad\psi(\gamma_{\max},\lambda)=\max_{\gamma\geq 0}\psi(\gamma,\lambda)=\frac{1}{\lambda}.$
  • (b)
    For a fixed $\lambda\in[1,+\infty)$, the function $\psi(\gamma,\lambda)$ is non-decreasing and upper-bounded. More precisely, $\psi(\gamma,\lambda)\leq 1,\quad\forall\gamma\geq 0.$
Proof.
First, note that for any fixed $\lambda$, we have
$\lim_{\gamma\rightarrow\infty}\psi(\gamma,\lambda)=1.$ (21)
Now, given $\lambda>0$, observe that
$\frac{\partial\psi}{\partial\gamma}(\gamma,\lambda)=\frac{2(\lambda-1)\gamma}{(\gamma^{2}+\lambda)^{2}}.$ (22)
  • (a)
    Consider $\lambda\in(0,1)$ fixed. As the maximizers of $\psi(\gamma,\lambda)$, for $\gamma>0$, must satisfy $\frac{\partial\psi}{\partial\gamma}(\gamma,\lambda)=0,$ from (22), we have $02(\lambda-1)\gamma=$ 0, thus $\gamma_{\max}=0$. Substituting $\gamma_{\max}$ into $\psi(\gamma,\lambda)$, we obtain
    $\psi(\gamma_{\max},\lambda)=1/2.$
  • (b)
    Now, for fixed $\lambda\in[1,+\infty)$, from (22), we have $\frac{\partial\psi}{\partial\gamma}(\gamma,\lambda)>0$ for $\gamma\geq 0$. Therefore, $\psi$ is increasing for $\gamma\in[0,+\infty)$, with $\psi(0,\lambda)=0$. Then, from (21), we conclude that $\psi(\gamma,\lambda)\leq 1$, for all $\gamma\geq 0$.
We are now ready to prove two key lemmas for the local convergence analysis. They show that the norm of LMMSS direction $d_{k}$ is ${\cal O}(\textup{dist}(x_{k},X^{*}))$, considering two distinct cases. Lemma 5 addresses the scenario where the rank of the Jacobian around the solution (stationary-point) is constant, while Lemma 6 deals with the case of diminishing rank.
Lemma 5.
Suppose that Assumptions 1-3 are valid in $B(x^{*},\delta)$, and
$rank(J(x)^{T}J(x))=rank(J(x^{*})^{T}J(x^{*}))=q\geq 1$
for all $x\in B(x^{*},\delta)$. If $x_{k}\in B(x^{*},\delta)$, $\lambda_{k}>0$ and $J_{k}^{T}F_{k}\neq 0$, then there exists $c_{1}>0$ such that
$\|d_{k}\|\leq c_{1}\textup{dist}(x_{k},X^{*}).$ (23)
Proof.
Recall that $J_{k}^{T}F_{k}\in\mathcal{R}(J_{k}^{T})=\mathcal{N}(J_{k})^{\perp}=\mathcal{N}(J_{k}^{T}J_{k})^{\perp}$, and
$(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)d_{k}=-J_{k}^{T}F_{k}.$ (24)
First, notice that $d_{k}\notin\mathcal{N}(J_{k})$. In fact, from (24), we have
$d_{k}^{T}(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)d_{k}=-(J_{k}d_{k})^{T}F_{k},$
which implies that $\left\lVert J_{k}d_{k}\right\rVert^{2}+\lambda_{k}\left\lVert Ld_{k}\right\rVert^{2}=-(J_{k}d_{k})^{T}F_{k}$, thus, if we assume that $d_{k}\in\mathcal{N}(J_{k})$, we conclude that $\lambda_{k}\left\lVert Ld_{k}\right\rVert^{2}=0$. But $\left\lVert Ld_{k}\right\rVert>0$, given that, by Assumption 1, $\mathcal{N}(J_{k})\cap\mathcal{N}(L)=0$, and also $\lambda_{k}\neq 0$, and we have a contradiction.
We can write $d_{k}=d_{N}+d_{R},$ with $d_{N}\in\mathcal{N}(J_{k})$ and $d_{R}\in\mathcal{N}(J_{k})^{\perp}=\mathcal{R}(J_{k}^{T})$. Since $d_{k}\notin\mathcal{N}(J_{k})$, we have $d_{R}\neq 0$.
Thus, from (24)
$-J_{k}^{T}F_{k}$ $=(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)d_{k}=(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)(d_{N}+d_{R})$ (25)
$=J_{k}^{T}J_{k}d_{N}+J_{k}^{T}J_{k}d_{R}+\lambda_{k}L^{T}Ld_{N}+\lambda_{k}L^{T}Ld_{R}$
$=J_{k}^{T}J_{k}d_{R}+\lambda_{k}L^{T}Ld_{N}+\lambda_{k}L^{T}Ld_{R}.$
When we multiply (25) by $d_{N}^{T}$, we get
$(J_{k}d_{N})^{T}J_{k}d_{R}+\lambda_{k}\left\lVert Ld_{N}\right\rVert^{2}+\lambda_{k}d_{N}^{T}L^{T}Ld_{R}=-(J_{k}d_{N})^{T}F_{k},$
and as $d_{N}\in\mathcal{N}(J_{k})$ and $\lambda_{k}>0$, we have:
$\left\lVert Ld_{N}\right\rVert^{2}=-(Ld_{N})^{T}Ld_{R}.$ (26)
From (26), we also conclude that $\left\lVert Ld_{N}\right\rVert^{2}\leq\left\lVert Ld_{N}\right\rVert\left\lVert Ld_{R}\right\rVert$, and as $d_{N}\notin{\cal N}(L)$ we obtain
$\left\lVert Ld_{N}\right\rVert\leq\left\lVert Ld_{R}\right\rVert.$ (27)
On the other hand, when we multiply (25) by $d_{R}^{T}$, we get
$\left\lVert J_{k}d_{R}\right\rVert^{2}+\lambda_{k}(Ld_{R})^{T}Ld_{N}+\lambda_{k}\left\lVert Ld_{R}\right\rVert^{2}=-(J_{k}d_{R})^{T}F_{k}.$ (28)
Using (26) in (28), we are left with
$\left\lVert J_{k}d_{R}\right\rVert^{2}+\lambda_{k}\left\lVert Ld_{R}\right\rVert^{2}=-(J_{k}d_{R})^{T}F_{k}+\lambda_{k}\left\lVert Ld_{N}\right\rVert^{2}.$
From this and (27), we have
$\left\lVert J_{k}d_{R}\right\rVert^{2}\leq-(J_{k}d_{R})^{T}F_{k}.$ (29)
Let $J_{k}=U\Sigma V^{T}$ be the SVD for $J_{k}$. As $d_{R}\in\mathcal{N}(J_{k})^{\perp}$, $d_{R}$ can be written as a linear combination of the columns of $V$ corresponding to the non-zero singular values of $J_{k}$. Therefore,
$\left\lVert J_{k}d_{R}\right\rVert=\left\lVert U\Sigma V^{T}d_{R}\right\rVert\geq s_{q,k}\left\lVert V^{T}d_{R}\right\rVert=s_{q,k}\left\lVert d_{R}\right\rVert,$
where $s_{q,k}$ is the smallest positive singular value of $J_{k}$. By using $s_{q,k}\left\lVert d_{R}\right\rVert\leq\left\lVert J_{k}d_{R}\right\rVert$ in (29):
$s_{q,k}^{2}\left\lVert d_{R}\right\rVert^{2}\leq-(J_{k}d_{R})^{T}F_{k}\leq\left\lVert J_{k}^{T}F_{k}\right\rVert\left\lVert d_{R}\right\rVert,$
which implies
$\left\lVert d_{R}\right\rVert\leq\frac{1}{s_{q,k}^{2}}\left\lVert J_{k}^{T}F_{k}\right\rVert.$ (30)
Now let us show that $d_{R}\notin\mathcal{N}(L)$. Suppose $d_{R}\in\mathcal{N}(L)$, from (26), we get $\left\lVert Ld_{N}\right\rVert^{2}=-d_{N}^{T}L^{T}Ld_{R}=0,$ implying that $d_{N}\in\mathcal{N}(L)$ and $d_{N}\in\mathcal{N}(J_{k})$, which is a contradiction with Assumption 1.
By Assumption 1, there exists $\gamma$ such that $\gamma\left\lVert d_{N}\right\rVert^{2}\leq\left\lVert J_{k}d_{N}\right\rVert^{2}+\left\lVert Ld_{N}\right\rVert^{2}$, which implies
$\left\lVert d_{N}\right\rVert^{2}\leq\frac{1}{\gamma}\left\lVert Ld_{N}\right\rVert^{2}.$ (31)
Therefore, from (27), (30), (31), and Assumption 2,
$\left\lVert d_{k}\right\rVert^{2}$ $=\left\lVert d_{N}+d_{R}\right\rVert^{2}=\left\lVert d_{N}\right\rVert^{2}+2d_{N}^{T}d_{R}+\left\lVert d_{R}\right\rVert^{2}=\left\lVert d_{N}\right\rVert^{2}+\left\lVert d_{R}\right\rVert^{2}$ (32)
$\leq\frac{1}{\gamma}\left\lVert Ld_{N}\right\rVert^{2}+\left\lVert d_{R}\right\rVert^{2}\leq\frac{1}{\gamma}\left\lVert Ld_{R}\right\rVert^{2}+\left\lVert d_{R}\right\rVert^{2}$
$\leq\Big{(}\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1\Big{)}\frac{1}{s_{q,k}^{4}}\left\lVert J_{k}^{T}F_{k}\right\rVert^{2}=\Big{(}\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1\Big{)}\frac{1}{s_{q,k}^{4}}\left\lVert J_{k}^{T}F_{k}-J_{*}^{T}F_{*}\right\rVert^{2}$
$\leq\Big{(}\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1\Big{)}\frac{1}{s_{q,k}^{4}}L_{3}^{2}\textup{dist}^{2}(x_{k},X^{*}).$
Since $J(x)$ is continuous, it is clear that for a sufficiently small $\delta>0$, condition (13) is satisfied for all $x_{k}\in B(x^{*},\delta)$. Finally, we have
$\left\lVert d_{k}\right\rVert\leq\Big{(}\sqrt{\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1}\Big{)}\frac{\kappa}{\lambda_{q}^{*}}L_{3}\textup{dist}(x_{k},X^{*}),$
where we recall that $\lambda_{q}^{*}$ is the smallest positive eigenvalue of $J_{*}^{T}J_{*}$. Then define $c_{1}=\Big{(}\sqrt{\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1}\Big{)}\frac{\kappa}{\lambda_{q}^{*}}L_{3}$ and we have the result. ∎
Lemma 6.
Suppose that Assumptions 13 are valid in $B(x^{*},\delta)$, for some $\delta>0$. If $x_{k}\in B(x^{*},\delta)$, $rank(J_{k})=\ell\geq rank(J_{*})=q\geq 1$, and Assumption 4 is satisfied, $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r},$ with $r\in(0,1]$, then there exists $c_{1}>0$ such that $\left\lVert d_{k}\right\rVert\leq c_{1}\textup{dist}(x_{k},X^{*})$.
Proof.
Consider the GSVD for the pair $(J_{k},L)$. In this case where $rank(J_{k})=\ell,rank(J_{*})=q\geq 1$ and $rank(L)=p$ with $n-q\leq p$ and $n-\ell\leq p$ (see Remark 1), we remark that:
  • For $1\leq i<n-\ell$, we have $\sigma_{i,k}$=0 and $\mu_{i,k}=1.$ This occurs because of the fact that $rank(J_{k})=\ell$ and the relationship between the singular values of the GSVD (see Theorem 1). Thus, in this case, $(\Gamma_{k})_{ii}=\frac{1}{\lambda_{k}}.$
  • For $n-\ell\leq i<n-q$, if $\sigma_{i,k}\rightarrow 0$, then $\mu_{i,k}\rightarrow 1$, and we have $(\Gamma_{k})_{ii}\rightarrow\frac{1}{\lambda_{k}}$.
  • For $n-q\leq i\leq p$, then $(\Gamma_{k})_{ii}=\frac{1}{\sigma_{i,k}+\lambda_{k}\mu_{i,k}}<\frac{1}{\sigma_{i,k}}\leq\frac{1}{\sigma^{*}_{\text{min}}},$ where $\sigma^{*}_{\text{min}}$ is the smallest positive singular value of $J(x^{*}).$
Note that, for $x_{k}\in B(x^{*},\delta)$ and due to the continuity of $\mu_{i,k}$, in the case where $n-\ell\leq i<n-q$ we can assume that $\mu_{i,k}>\frac{1}{2}$ and thus, for $\delta$ sufficiently small, we have $(\Gamma_{k})_{ii}\leq\frac{1}{4\lambda_{k}}$. Therefore,
$\left\lVert\left[\begin{array}[]{cc}\Gamma_{k}&0\\ 0&I_{n-p}\end{array}\right]\right\rVert\leq\max{\Big{[}\frac{1}{\lambda_{k}},\frac{1}{4\lambda_{k}},\frac{1}{\sigma^{*}_{\text{min}}},1\Big{]}}\leq\frac{1}{\lambda_{k}},$ (33)
as $x_{k}\in B(x^{*},\delta)$ and $\left\lVert J_{k}^{T}F_{k}\right\rVert<1.$
From (19) and (33), we obtain
$\left\lVert d_{k}\right\rVert$ $=\left\lVert X_{k}\left[\begin{array}[]{cc}\Gamma_{k}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{T}{J_{k}}^{T}F_{k}\right\rVert$ (34)
$=\left\lVert X_{k}\left[\begin{array}[]{cc}\Gamma_{k}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{T}\Big{(}{J_{k}}^{T}F_{k}-J(\bar{x}_{k})^{T}F(\bar{x}_{k})\Big{)}\right\rVert$
$\leq\left\lVert X_{k}\right\rVert^{2}\frac{1}{\lambda_{k}}\left\lVert{J_{k}}^{T}F_{k}-J(\bar{x}_{k})^{T}F(\bar{x}_{k})-J_{k}^{T}J_{k}(x_{k}-\bar{x}_{k})\right\rVert$
$+\left\lVert X_{k}\left[\begin{array}[]{cc}\Gamma_{k}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{T}J_{k}^{T}J_{k}(x_{k}-\bar{x}_{k})\right\rVert.$
Observe from GSVD that
$X_{k}J_{k}^{T}J_{k}$ $=X_{k}X_{k}^{-T}\left[\begin{array}[]{cc}\Sigma_{k}&0\\ 0&I_{n-p}\end{array}\right]U_{k}^{T}U_{k}\left[\begin{array}[]{cc}\Sigma_{k}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{-1}=\left[\begin{array}[]{cc}\Sigma_{k}^{2}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{-1},$
then, as $\Gamma_{k}:=(\Sigma_{k}^{2}+\lambda_{k}M_{k}^{2})^{-1}$,
$\left\lVert X_{k}\left[\begin{array}[]{cc}\Gamma_{k}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{T}J_{k}^{T}J_{k}(x_{k}-\bar{x}_{k})\right\rVert$ $\leq\left\lVert X_{k}\right\rVert\zeta\left\lVert X_{k}^{-1}\right\rVert\left\lVert x_{k}-\bar{x}_{k}\right\rVert$ (35)
$\leq\left\lVert X_{k}\right\rVert\left\lVert X_{k}^{-1}\right\rVert\left\lVert x_{k}-\bar{x}_{k}\right\rVert,$
where this last inequality comes from the fact that
$\left\lVert\left[\begin{array}[]{cc}(\Sigma_{k}^{2}+\lambda_{k}M_{k}^{2})^{-1}\Sigma_{k}^{2}&0\\ 0&I_{n-p}\end{array}\right]\right\rVert\leq\max\Big{\{}\max_{1\leq i\leq p}{\Big{\{}\frac{\sigma_{i,k}^{2}}{\sigma_{i,k}^{2}+\lambda_{k}\mu_{i,k}^{2}}}\Big{\}},1\Big{\}}=:\zeta\leq 1$
with $\sigma_{i,k}$ and $\mu_{i,k}$ as in (17).
Therefore, using (19), (35), (9), (20) and Lemma 3 in (34), we have
$\left\lVert d_{k}\right\rVert$ $\leq\left\lVert X_{k}\right\rVert^{2}\frac{1}{\lambda_{k}}\left\lVert{J_{k}}^{T}F_{k}-J(\bar{x}_{k})^{T}F(\bar{x}_{k})-J_{k}^{T}J_{k}(x_{k}-\bar{x}_{k})\right\rVert$ (36)
$+\left\lVert X_{k}\right\rVert\left\lVert X_{k}^{-1}\right\rVert\left\lVert x_{k}-\bar{x}_{k}\right\rVert$
$\leq\frac{1}{\gamma\lambda_{k}}L_{4}\left\lVert x_{k}-\bar{x}_{k}\right\rVert^{2}+\frac{1}{\gamma\lambda_{k}}\left\lVert J_{k}^{T}F(\bar{x}_{k})\right\rVert+\frac{\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}}{\sqrt{\gamma}}\left\lVert x_{k}-\bar{x}_{k}\right\rVert.$
Finally, from Assumption 4 and $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$, the result follows. ∎
4 Local Convergence
Lemmas 5 and 6 showed that $\left\lVert d_{k}\right\rVert\leq c_{1}\left\lVert x_{k}-\bar{x}_{k}\right\rVert$, under specific assumptions. As we shall see in this section, such inequality is key for the local convergence analysis under an error bound condition (Assumption 3).
Apart from the next lemma which is an intermediate result, the remaining of this section is organized in subsections according to whether the Jacobian rank near the solution set is constant or not.
Throughout this section, we consider the LMMSS iteration (2)–(3) with $\alpha_{k}=1$ for every $k$, that is, the “pure” LMMSS.
Lemma 7.
Let $\{x_{k}\}$ be the sequence generated by the LMMSS (with $\alpha_{k}=1$) and suppose Assumptions 1-3 hold. If $x_{k+1},x_{k}\in B(x^{*},\delta)$ and $\left\lVert d_{k}\right\rVert\leq c_{1}\left\lVert x_{k}-\bar{x}_{k}\right\rVert$, then
$\omega\textup{dist}(x_{k+1},X^{*})$ $\leq(L_{4}c_{1}^{2}+L_{5})||x_{k}-\bar{x}_{k}||^{2}+\lambda_{k}\left\lVert L\right\rVert^{2}c_{1}||x_{k}-\bar{x}_{k}||$ (37)
$+||J(x_{k})^{T}F(\bar{x}_{k})||+||J(x_{k+1})^{T}F(\bar{x}_{k})||,$
where $L_{5}:=L_{0}L_{2}(1+c_{1})(2+c_{1})$.
Proof.
For any $x,y\in\mathbb{R}^{n}$, the reverse triangle inequality implies
$\|\nabla\phi(y)-\nabla\phi(x)-J(x)^{T}J(x)(y-x)\|\geq\|\nabla\phi(y)\|-\|\nabla\phi(x)-J(x)^{T}J(x)(y-x)\|.$
Thus, for $y=x_{k+1}$, $x=x_{k}$, $x_{k+1},x_{k}\in B(x^{*},\delta)$ and using Lemma 1, we have:
$\|\nabla\phi(x_{k+1})\|$ $\leq\|\nabla\phi(x_{k+1})-\nabla\phi(x_{k})-J_{k}^{T}J_{k}d_{k}\|+\|\nabla\phi(x_{k})+J_{k}^{T}J_{k}d_{k}\|$ (38)
$\leq L_{4}\|d_{k}\|^{2}+\|(J_{k}-J_{k+1})^{T}F_{k+1}\|+\|J_{k}^{T}F_{k}+J_{k}^{T}J_{k}d_{k}\|.$
From the error bound condition (Assumption 3), the definition of the LMMSS iteration and (38), we get
$\omega\textup{dist}(x_{k+1},X^{*})\leq L_{4}\|d_{k}\|^{2}+\|(J_{k}-J_{k+1})^{T}F_{k+1}\|+\lambda_{k}\left\lVert L\right\rVert^{2}\left\lVert d_{k}\right\rVert.$ (39)
Now, rewriting (8), we have
$\|(J_{k}-J_{k+1})^{T}F_{k+1}\|$ $\leq L_{0}L_{2}\|x_{k}-\bar{x}_{k}\|\|x_{k+1}-\bar{x}_{k}\|+\|J_{k}^{T}F(\bar{x}_{k})\|$ (40)
$+L_{0}L_{2}\|x_{k+1}-\bar{x}_{k}\|^{2}+\|J_{k+1}^{T}F(\bar{x}_{k})\|.$
Furthermore, since $\left\lVert d_{k}\right\rVert\leq c_{1}\left\lVert x_{k}-\bar{x}_{k}\right\rVert$, it holds
$\|x_{k+1}-\bar{x}_{k}\|=\|x_{k}-\bar{x}_{k}+x_{k+1}-x_{k}\|\leq\|x_{k}-\bar{x}_{k}\|+\|d_{k}\|\leq(1+c_{1})\|x_{k}-\bar{x}_{k}\|.$ (41)
Thus, using (40) and (41) in (39) yields
$\omega\textup{dist}(x_{k+1},X^{*})$ $\leq L_{4}\|d_{k}\|^{2}+L_{0}L_{2}\|x_{k}-\bar{x}_{k}\|\|x_{k+1}-\bar{x}_{k}\|+\|J_{k}^{T}F(\bar{x}_{k})\|$ (42)
$+L_{0}L_{2}\|x_{k+1}-\bar{x}_{k}\|^{2}+\|J_{k+1}F(\bar{x}_{k})\|+\lambda_{k}\left\lVert L\right\rVert^{2}\left\lVert d_{k}\right\rVert$
$\leq L_{4}\|d_{k}\|^{2}+L_{0}L_{2}(1+c_{1})\|x_{k}-\bar{x}_{k}\|^{2}+\|J_{k}^{T}F(\bar{x}_{k})\|$
$+L_{0}L_{2}(1+c_{1})^{2}\|x_{k}-\bar{x}_{k}\|^{2}+\|J_{k+1}^{T}F(\bar{x}_{k})\|+\lambda_{k}\left\lVert L\right\rVert^{2}\left\lVert d_{k}\right\rVert$
$\leq L_{4}c_{1}^{2}\|x_{k}-\bar{x}_{k}\|^{2}+\lambda_{k}\left\lVert L\right\rVert^{2}c_{1}\|x_{k}-\bar{x}_{k}\|$
$+L_{0}L_{2}(1+c_{1})\|x_{k}-\bar{x}_{k}\|^{2}+L_{0}L_{2}(1+c_{1})^{2}\|x_{k}-\bar{x}_{k}\|^{2}$
$+\|J_{k}^{T}F(\bar{x}_{k})\|+\|J_{k+1}^{T}F(\bar{x}_{k})\|.$
4.1 Constant Rank
In this section, we assume that $1\leq\text{rank}(J(x^{*}))=q\leq\min\{n,m\}$, and $\text{rank}(J(x))=\text{rank}(J(x^{*}))$ for each $x\in B(x^{*},\delta)$. From Lemma 5, we recall that $\left\|d_{k}\right\|\leq c_{1}\text{dist}(x_{k},X^{*})$, with $c_{1}=\Big{(}\sqrt{\frac{\left\lVert L\right\rVert^{2}}{\gamma}+1}\Big{)}\frac{\kappa}{\lambda_{p}^{*}}L_{3}$, provided that $\lambda_{k}>0,$ for $x_{k}\neq x^{*}$.
Next, we present two lemmas that will aid in proving the convergence theorem under Assumption 4.
Lemma 8.
Suppose that Assumptions 1-3 are valid, and $rank(J(x))=rank(J(x^{*}))\geq 1$ in $B(x^{*},\delta)$. Additionally, assume that Assumption 4 is verified, and $\lambda_{k}=\|J_{k}^{T}F_{k}\|$. If $x_{k},x_{k+1}\in B(x^{*},\delta/2)$ and $\textup{dist}(x_{k},X^{*})<\varepsilon$, with
$\varepsilon\leq\left(\frac{\eta\omega}{\tilde{C}}\right)^{1/r},$ (43)
with $\eta\in(0,1)$ and $L_{6}=L_{4}c_{1}^{2}+L_{5}+L_{3}\left\lVert L\right\rVert^{2}c_{1}$, then
$\textup{dist}(x_{k+1},X^{*})\leq\eta\,\textup{dist}(x_{k},X^{*}).$
Proof.
We know that $J(x)$ has constant rank in $B(x^{*},\delta)$, $\lambda_{k}=\|J_{k}^{T}F_{k}\|$, and Assumptions 4 is satisfied. Thus, from (37) and (7), we have
$\omega\textup{dist}(x_{k+1},X^{*})$ $\leq(L_{4}c_{1}^{2}+L_{5})||x_{k}-\bar{x}_{k}||^{2}+\lambda_{k}\left\lVert L\right\rVert^{2}c_{1}||x_{k}-\bar{x}_{k}||$ (44)
$+||J(x_{k})^{T}F(\bar{x}_{k})||+||J(x_{k+1})^{T}F(\bar{x}_{k})||$
$\leq(L_{4}c_{1}^{2}+L_{5})\|x_{k}-\bar{x}_{k}\|^{2}+L_{3}\left\lVert L\right\rVert^{2}c_{1}\|x_{k}-\bar{x}_{k}\|^{2}$
$+C\|x_{k}-\bar{x}_{k}\|^{1+r}+C\|x_{k+1}-\bar{x}_{k}\|^{1+r}$
$\leq(L_{4}c_{1}^{2}+L_{5}+L_{3}\left\lVert L\right\rVert^{2}c_{1})\|x_{k}-\bar{x}_{k}\|^{2}+C\|x_{k}-\bar{x}_{k}\|^{1+r}$
$+C(1+c_{1})^{1+r}\|x_{k}-\bar{x}_{k}\|^{1+r}$
$=L_{6}\|x_{k}-\bar{x}_{k}\|^{2}+(1+(1+c_{1})^{1+r})C\|x_{k}-\bar{x}_{k}\|^{1+r}$
$\leq L_{6}\|x_{k}-\bar{x}_{k}\|^{1+r}+(1+(1+c_{1})^{1+r})C\|x_{k}-\bar{x}_{k}\|^{1+r}$
$=\Big{[}L_{6}+(1+(1+c_{1})^{1+r})C\Big{]}\left\lVert x_{k}-\bar{x}_{k}\right\rVert^{1+r}$
$=:\tilde{C}\left\lVert x_{k}-\bar{x}_{k}\right\rVert^{1+r}=\tilde{C}\textup{dist}(x_{k},X^{*})^{1+r}$
where $L_{6}:=L_{4}c_{1}^{2}+L_{5}+L_{3}\left\lVert L\right\rVert^{2}c_{1}$ and $\tilde{C}:=L_{6}+(1+(1+c_{1})^{1+r})C$.
From (44), we have
$\textup{dist}(x_{k+1},X^{*})\leq\dfrac{\tilde{C}}{\omega}\textup{dist}(x_{k},X^{*})^{1+r}\leq\dfrac{\tilde{C}}{\omega}\varepsilon^{r}\textup{dist}(x_{k},X^{*})$
and, given $\eta\in]0,1[$, for $\varepsilon\leq\left(\frac{\eta\omega}{\tilde{C}}\right)^{1/r}$ we obtain
$\textup{dist}(x_{k+1},X^{*})\leq\eta\textup{dist}(x_{k},X^{*}).$
Lemma 9.
Suppose that the assumptions of Lemma 8 are satisfied, and
$\varepsilon=\min\left\{\frac{\frac{\delta}{2}}{1+\frac{c_{1}}{1-\eta}},\left(\frac{\eta\omega}{\tilde{C}}\right)^{1/r}\right\}.$ (45)
If $x_{0}\in B(x^{*},\varepsilon)$, then $x_{k+1}\in B(x^{*},\frac{\delta}{2})$ and $\textup{dist}(x_{k},X^{*})\leq\varepsilon$, for every $k\in\mathbb{N}$.
Proof.
The proof is done by induction. For $k=0$, we have $\textup{dist}(x_{0},X^{*})\leq\|x_{0}-x^{*}\|\leq\varepsilon$ and
$\|x_{1}-x^{*}\|\leq\|x_{1}-x_{0}\|+\|x_{0}-x^{*}\|\leq\|d_{0}\|+\varepsilon\leq c_{1}\textup{dist}(x_{0},X^{*})+\varepsilon\leq(1+c_{1})\varepsilon<\frac{\delta}{2}.$
For $k\geq 1$, assume that $x_{i}\in B(x^{*},\frac{\delta}{2})$ and $\textup{dist}(x_{i-1},X^{*})\leq\varepsilon$ for $i=1,...,k$. Then, from Lemma 8 applied to $x_{k-1}$ and $x_{k}$, we obtain
$\|x_{k}-\bar{x}_{k}\|=\textup{dist}(x_{k},X^{*})\leq\eta\textup{dist}(x_{k-1},X^{*})\leq\eta\varepsilon<\varepsilon.$
Moreover, by Lemma 8 and the induction hypothesis for $i\leq k$, we have
$\textup{dist}(x_{i},X^{*})\leq\eta\textup{dist}(x_{i-1},X^{*})\leq\cdots\leq\eta^{i}\textup{dist}(x_{0},X^{*})\leq\eta^{i}\varepsilon<\varepsilon.$
Thus, using Lemma 5,
$\|x_{k+1}-x^{*}\|$ $\leq\|x_{1}-x^{*}\|+\sum_{i=1}^{k}\|d_{i}\|\leq\|x_{1}-x^{*}\|+\sum_{i=1}^{k}c_{1}\textup{dist}(x_{i},X^{*})$ (46)
$\leq(1+c_{1})\varepsilon+c_{1}\varepsilon\sum_{i=1}^{\infty}\eta^{i}=\Big{(}1+\frac{c_{1}}{1-\eta}\Big{)}\varepsilon\leq\frac{\delta}{2}.$
Theorem 2.
Suppose that Assumptions 1-3 are valid, and $rank(J(x))=rank(J(x^{*}))\geq 1$ in $B(x^{*},\delta)$. Let $\{x_{k}\}$ be generated by the LMMSS method with $\alpha_{k}=1$, $\lambda_{k}=\|J_{k}^{T}F_{k}\|$ for all $k$. If Assumption 4 is verified and $x_{0}\in B(x^{*},\varepsilon)$, where $\varepsilon>0$ is given by (45), then $\{\textup{dist}(x_{k},X^{*})\}$ converges superlinearly to zero. Moreover, the sequence $\{x_{k}\}$ converges to a solution $\bar{x}\in X^{*}\cap B(x^{*},\frac{\delta}{2})$.
Proof.
From (44) we obtain
$\dfrac{\textup{dist}(x_{k+1},X^{*})}{\textup{dist}(x_{k},X^{*})}\leq\dfrac{\tilde{C}}{\omega}\textup{dist}(x_{k},X^{*})^{r}$
and since $\textup{dist}(x_{k},X^{*})^{r}\rightarrow 0$ (due to Lemmas 8 and 9) we have that $\{\textup{dist}(x_{k},X^{*})\}$ converges to zero superlinearly.
Once $\textup{dist}(x_{k},X^{*})$ converges to zero and $x_{k}\in B(x^{*},\frac{\delta}{2})$ for every $k$, it remains to show that $\{x_{k}\}$ converges. From Lemma 5 and Lemma 8, we have
$\|d_{k}\|\leq c_{1}\text{dist}(x_{k},X^{*})\leq c_{1}\eta^{k}\text{dist}(x_{0},X^{*})\leq c_{1}\varepsilon\eta^{k},$
for every $k\geq 1$. Thus, for any positive integers $\ell$ and $q$, with $\ell\geq q$,
$\|x_{\ell}-x_{q}\|\leq\sum_{i=q}^{\ell-1}\|d_{i}\|\leq\sum_{i=q}^{\infty}\|d_{i}\|\leq c_{1}\varepsilon\sum_{i=q}^{\infty}\eta^{i},$
implying that $\{x_{k}\}\subset\mathbb{R}^{n}$ is a Cauchy sequence and, therefore, converges. ∎∎
Remark 3.
In particular, when $r=1$ in Assumption 4, $\textup{dist}(x_{k},X^{*})$ converges to zero quadratically.
4.2 Diminishing Rank
Since the matrix rank is a lower semicontinuous function, for a given sequence $\{x_{k}\}$ converging to $x^{*}\in X^{*}$ it may happen that $\text{rank}J(x_{k})>\text{rank}J(x^{*})$, for every $k$ sufficiently large. We refer to this situation as “diminishing rank” scenario. In this case the convergence results will depend on the specific choice $\lambda_{k}=\|J(x_{k})^{T}F(x_{k})\|^{r^{\prime}}$, with $r^{\prime}\leq r$ from Assumption 4. To simplify the exposition, we consider $r^{\prime}=r$ in the results that follow.
Lemma 10.
Suppose that Assumptions 1-3 and 4 are valid, and $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$. If $x_{k},x_{k+1}\in B(x^{*},\delta/2)$ and $\textup{dist}(x_{k},X^{*})<\varepsilon$, where
$\varepsilon\leq\Big{(}\frac{\eta\omega}{\hat{C}}\Big{)}^{1/r}$ (47)
with $\eta\in(0,1)$, then
$\textup{dist}(x_{k+1},X^{*})\leq\eta\,\textup{dist}(x_{k},X^{*}).$
Proof.
Note that from Assumption 4 and Lemma 6, for $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r}$, we have $\left\lVert d_{k}\right\rVert\leq c_{1}\textup{dist}(x_{k},X^{*})$, with
$c_{1}=\frac{1}{\gamma\omega^{r}}(L_{4}+C)+\frac{\sqrt{L_{2}^{2}+\left\lVert L\right\rVert^{2}}}{\sqrt{\gamma}}.$ (48)
From Lemma 7, Assumption 4, $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r}$, (10), $\textup{dist}(x_{k},X^{*})\leq\delta/2<1/2$ and assuming that $\textup{dist}(x_{k},X^{*})<\varepsilon$, we have
$\omega\textup{dist}(x_{k+1},X^{*})$ $\leq(L_{4}c_{1}^{2}+L_{5})||x_{k}-\bar{x}_{k}||^{2}+\lambda_{k}\left\lVert L\right\rVert^{2}c_{1}||x_{k}-\bar{x}_{k}||+$ (49)
$+||J(x_{k})^{T}F(\bar{x}_{k})||+||J(x_{k+1})^{T}F(\bar{x}_{k})||$
$\leq(L_{4}c_{1}^{2}+L_{5})\|x_{k}-\bar{x}_{k}\|^{2}+L_{3}^{r}\left\lVert L\right\rVert^{2}c_{1}\|x_{k}-\bar{x}_{k}\|^{1+r}$
$+C\|x_{k}-\bar{x}_{k}\|^{1+r}+C\|x_{k+1}-\bar{x}_{k}\|^{1+r}$
$\leq(L_{4}c_{1}^{2}+L_{5}+L_{3}^{r}\left\lVert L\right\rVert^{2}c_{1})\|x_{k}-\bar{x}_{k}\|^{2}$
$+(1+(1+c_{1})^{1+r})C\|x_{k}-\bar{x}_{k}\|^{1+r}$
$=[L_{7}+(1+(1+c_{1})^{1+r})C]\|x_{k}-\bar{x}_{k}\|^{1+r}$
$\leq[L_{7}+(1+(1+c_{1})^{1+r})C]\varepsilon^{r}\left\lVert x_{k}-\bar{x}_{k}\right\rVert$
$:=\hat{C}\varepsilon^{r}\left\lVert x_{k}-\bar{x}_{k}\right\rVert,$
with $c_{1}$ as in (48) and $L_{7}=L_{4}c_{1}^{2}+L_{5}+L_{3}^{r}\left\lVert L\right\rVert^{2}c_{1}$.
Therefore, for $\varepsilon\leq\Big{(}\frac{\eta\omega}{\hat{C}}\Big{)}^{1/r},$ we have $\textup{dist}(x_{k+1},X^{*})\leq\eta\,\textup{dist}(x_{k},X^{*})$. ∎∎
Using an appropriate $\varepsilon>0$, and Lemma 10 , the proofs of the next results are analogous to those of Lemma 9 and Theorem 2, respectively.
Lemma 11.
Suppose that the Assumptions of Lemma 10 are satisfied, and consider $\varepsilon$ as
$\varepsilon=\min\left\{\frac{\frac{\delta}{2}}{1+\frac{c_{1}}{1-\eta}},\Big{(}\frac{\eta\omega}{\hat{C}}\Big{)}^{1/r}\right\}.$ (50)
If $x_{0}\in B(x^{*},\varepsilon)$, then $x_{k+1}\in B(x^{*},\frac{\delta}{2})$, and $\textup{dist}(x_{k},X^{*})\leq\varepsilon$ for every $k\in\mathbb{N}$.
Theorem 3.
Suppose that Assumptions 1-3 and 4 are valid. Let $\{x_{k}\}$ be generated by LMMSS with $\alpha_{k}=1$, $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$, for all $k$. If $x_{0}\in B(x^{*},\varepsilon)$, where $\varepsilon>0$ as in (50), then $\{dist(x_{k},X^{*})\}$ converges superlinearly to zero. Furthermore, the sequence $\{x_{k}\}$ converges to the solution $\bar{x}\in X^{*}\cap B(x^{*},\delta/2).$
Remark 4.
The main difference from Theorem 2 is that when Assumption 4 holds with $10<r<$ 1 we cannot choose $\lambda_{k}=\|J_{k}^{T}F_{k}\|$ anymore; $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r}$ is essential for the convergence proof in this case. This becomes clear from the proof of Lemma 6; see inequality (36). In fact, we could choose $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r^{\prime}}$ with $r^{\prime}\leq r$ that all analysis remains valid with minor modifications.
5 Global Convergence
Now we consider Algorithm 1, a version of LMMSS globalized with line-search for non-zero residue nonlinear least-squares. It is worth to point out that Algorithm 1 differs from the algorithm proposed in [5] in the choice of the LM parameter $\lambda_{k}$ as well as the full step ($\alpha_{k}=1$) acceptance criterion.
Input: $\nu,\zeta,\vartheta,\theta\in(0,1)$, $M>0$ $F$, $J$, $L$, and $x_{0}\in\Omega$.
Set $\lambda_{0}=\left\lVert\nabla\phi({x}_{0})\right\rVert$, $k=0$.
  1. 1.
    If $\|\nabla\phi(x_{k})\|=0$, stop and return $\bar{x}=x_{k}$.
  2. 2.
    Compute $d_{k}$, a solution of $(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)d=-J_{k}^{T}F_{k}$.
  3. 3.
    If $\left\lVert\nabla\phi(x_{k}+d_{k})\right\rVert\leq\vartheta\left\lVert\nabla\phi(x_{k})\right\rVert$, set $\alpha_{k}=1$ and go to Step 6.
  4. 4.
    If $\|d_{k}\|>M$ or $-\nabla\phi(x_{k})^{T}d_{k}<\theta\|\nabla\phi(x_{k})\|^{2}$, set
    $d_{k}=-(J_{k}^{T}J_{k}+\lambda_{k}I)^{-1}J_{k}^{T}F_{k}.$
  5. 5.
    Choose $m$ as the smallest non-negative integer such that
    $\phi(x_{k}+\zeta^{m}{d}_{k})-\phi({x}_{k})\leq\nu\zeta^{m}\nabla\phi({x}_{k})^{T}{d}_{k}$
    Set $\alpha_{k}=\zeta^{m}$.
  6. 6.
    Update ${x}_{k+1}={x}_{k}+\alpha_{k}{d}_{k}$, $\lambda_{k+1}=\left\lVert\nabla\phi(x_{k+1})\right\rVert$, $k\leftarrow k+1$,
and go to Step 1.
Algorithm 1 LMMSS with line-search
Initially, we will demonstrate that the sequence of directions $d_{k}$ generated by Algorithm 1 is gradient-related (see [4, Sec. 1.2]). Subsequently, we will establish that any limit point of the sequence produced by this algorithm is a stationary point for (1), regardless of the initial point.
In this section, we assume that Assumption 4 is satisfied with $r=1$ to simplify notation and proofs.
We recall the definition of gradient-related directions from [4, Sec. 1.2].
Definition 1 (Gradient-related).
Let $\{x_{k}\}$ and $\{d_{k}\}$ be sequences in $\mathbb{R}^{n}$. The sequence $\{d_{k}\}$ is said to be gradient-related to $\{x_{k}\}$ if, for each subsequence $\{x_{k}\}_{k\in\mathcal{K}}$ (with $\mathcal{K}\subseteq\mathbb{N}$) converging to a non-stationary point of $\phi$, the corresponding subsequence $\{d_{k}\}_{k\in\mathcal{K}}$ is bounded and satisfies
$\limsup_{k\rightarrow\infty,\ k\in\mathcal{K}}\nabla\phi(x_{k})^{T}d_{k}<0.$ (51)
Proposition 1.
Let $\{d_{k}\}$, $\{x_{k}\}$ be sequences generated by Algorithm 1 starting from $x_{0}\in\mathbb{R}^{n}$. Suppose that Assumption 1 holds in the level set ${\cal L}_{0}=\{x\in\mathbb{R}^{n}\mid\phi(x)\leq\phi(x_{0})\}$. Then, $\{d_{k}\}$ is gradient-related.
Proof.
Suppose that Algorithm 1 generates an infinite sequence $\{x_{k}\}$, i.e., $\|\nabla\phi(x_{k})\|=\|J_{k}^{T}F_{k}\|>0$, for each $k$, then, consequently, $\left\lVert F(x_{k})\right\rVert>0$, and let $\{x_{k}\}_{k\in\mathcal{K}}$, $\mathcal{K}\subseteq\mathbb{N}$, be a subsequence converging to $x_{\infty}$, a non-stationary point for $\phi$, i.e., $\nabla\phi(x_{\infty})\neq 0$. From now on, consider $k\in\mathcal{K}$. From (18) and the definition of $d_{k}$, it follows that
$\nabla\phi(x_{k})^{T}d_{k}$ $=-d_{k}^{T}(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)d_{k}$
$=-d_{k}^{T}X_{k}^{-T}\left[\begin{array}[]{cc}\Sigma_{k}^{2}+\lambda_{k}M_{k}^{2}&0\\ 0&I_{n-p}\end{array}\right]X_{k}^{-1}d_{k}$
$:=-d_{k}^{T}X_{k}^{-T}\Lambda_{k}X_{k}^{-1}d_{k}.$
By Ostrowski’s inertia law, we have
$\lambda_{n}(X_{k}^{-T}\Lambda_{k}X_{k}^{-1})=\sigma_{n}^{2}(X_{k}^{-1})\lambda_{n}(\Lambda_{k}),$
where, with an abuse of notation, $\lambda_{n}(B)$ and $\sigma_{n}(B)$ represent the smallest eigenvalue and the smallest singular value of a matrix $B$, respectively. Then, from Lemma 3 it follows that
$\lambda_{n}(X_{k}^{-T}\Lambda_{k}X_{k}^{-1})\geq\gamma\lambda_{n}(\Lambda_{k})\geq\gamma\min\{1,\lambda_{k}\},$
where, in the last inequality, we use the fact that $\sigma_{i,k}^{2}+\mu_{i,k}^{2}=1$. Since $x_{\infty}$ is non-stationary, and $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert$, there exists a $\underline{\lambda}>0$ such that for $k\in{\cal K}$, $\lambda_{k}\geq\underline{\lambda}$. Thus, we obtain
$\nabla\phi(x_{k})^{T}d_{k}\leq-\gamma\min\{1,\underline{\lambda}\}\|d_{k}\|^{2}.$ (52)
On the other hand,
$\|\nabla\phi(x_{k})\|$ $=\|(J_{k}^{T}J_{k}+\lambda_{k}L^{T}L)d_{k}\|\leq\|J_{k}^{T}J_{k}+\lambda_{k}L^{T}L\|\|d_{k}\|$
$\leq\left(\|J_{k}\|^{2}+\lambda_{k}\|L\|^{2}\right)\|d_{k}\|.$
By the continuity of $F(x)$ and $J(x)$ and the convergence of the subsequence $\{x_{k}\}_{\mathcal{K}}$, there exist $L_{J}>0$ and $M_{0}>0$ such that $\|J(x_{k})^{T}F(x_{k})\|\leq M_{0}$ and $\left\lVert J(x_{k})\right\rVert\leq L_{J}$ for each $k$.
Thus, since $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert$, we have
$\|\nabla\phi(x_{k})\|\leq\left(L_{J}^{2}+M_{0}\|L\|^{2}\right)\|d_{k}\|=:M_{1}\|d_{k}\|,\quad\forall k\in{\cal K}.$ (53)
From (53) and (52), we obtain
$\nabla\phi(x_{k})^{T}d_{k}\leq-\frac{\gamma\min\{1,\underline{\lambda}\}}{M_{1}^{2}}\|\nabla\phi(x_{k})\|^{2},\quad\forall k\in{\cal K},$
which implies that
$\limsup_{k\rightarrow\infty,\ k\in\mathcal{K}}\nabla\phi(x_{k})^{T}d_{k}\leq-\frac{\gamma\min\{1,\underline{\lambda}\}}{M_{1}^{2}}\|\nabla\phi(x_{\infty})\|^{2}<0.$
Now we just need to show that $\{d_{k}\}_{k\in\mathcal{K}}$ is bounded. Recall from (19) that
$\|d_{k}\|\leq\|X_{k}\|^{2}\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|,\quad\forall k\geq 0,$ (54)
with $\Gamma_{k}=(\Sigma_{k}^{2}+\lambda_{k}M_{k}^{2})^{-1}$.
From Lemma 3, we have $\|X_{k}\|\leq 1/\sqrt{\gamma}$. Now let us analyze the term $\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|$. From Lemma 4, it follows that
$\|\Gamma_{k}\|\leq\dfrac{1}{\lambda_{k}},\quad\text{if}\quad 0<\lambda_{k}<1,\quad\text{and}\quad\|\Gamma_{k}\|\leq 1,\quad\text{if}\quad\lambda_{k}\geq 1.$ (55)
We separate into two cases to analyze such upper bound.
  • (a)
    If $10<\lambda_{k}<$ 1, then $\max\{\|\Gamma_{k}\|,1\}\leq\dfrac{1}{\lambda_{k}}$. Since $\lambda_{k}=\|J_{k}^{T}F_{k}\|$, we have
    $\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|\leq\dfrac{1}{\lambda_{k}}\|J_{k}^{T}F_{k}\|=\dfrac{1}{\|J_{k}^{T}F_{k}\|}\|J_{k}^{T}F_{k}\|=1.$
  • (b)
    If $\lambda_{k}\geq 1$, then $\max\{\|\Gamma_{k}\|,1\}\leq 1$. Thus, $\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|\leq M_{0}$.
Therefore, from (a) and (b), we conclude that
$\max\{\|\Gamma_{k}\|,1\}\|J_{k}^{T}F_{k}\|\leq\max\{1,M_{0}\},\quad\forall k\in{\cal K}.$
and
$\|d_{k}\|\leq\frac{1}{\gamma}\max\{1,M_{0}\}=:M_{2},\quad\forall k\in{\cal K}.$ (56)
Inequalities (52) and (56) show that for $k\in{\cal K}$, $d_{k}$ is the one determined in Step 2 of Algorithm 1 because it passes the seize of Step 4 for $M=M_{2}$ and $\theta=\gamma\min\{1,\underline{\lambda}\}/M_{1}^{2}$. Thus, the proof is complete. ∎∎
Remark 5.
The above proposition shows that LMMSS directions generated by Algorithm 1 are gradient-related as long as the completeness condition (Assumption 1) holds at every iterate – which also ensures that the solution of (2) is unique. Nevertheless, apart from special cases, e.g when the Jacobian $J(x)$ is full-rank everywhere, we cannot ensure that Assumption 1 holds at every $x_{k}$ generated by Algorithm 1. Hence, we safeguard the algorithm by using the classic LMM direction whenever the LMMSS direction fails to fulfill sufficient conditions for gradient-related property, namely, $\|d_{k}\|\leq M$, for $M>0$ and $-\nabla\phi(x_{k})^{T}d_{k}\geq\theta\|\nabla\phi(x_{k})\|^{p_{1}}$, for $\theta\in(0,1)$ and $p_{1}>0$.
If such conditions are not satisfied, we take $d_{k}=-(J_{k}^{T}J_{k}+\lambda_{k}I)^{-1}J_{k}^{T}F_{k}$ which yields gradient-related directions. In fact, from
$(J_{k}^{T}J_{k}+\lambda_{k}I)d_{k}=-\nabla\phi(x_{k}),$
we obtain $-\nabla\phi(x_{k})^{T}d_{k}=\|J_{k}d_{k}\|^{2}+\lambda_{k}\|d_{k}\|^{2}\geq\lambda_{k}\|d_{k}\|^{2},$ and since
$\|\nabla\phi(x_{k})\|=\|(J_{k}^{T}J_{k}+\lambda_{k}I)d_{k}\|\leq\|J_{k}^{T}J_{k}+\lambda_{k}I\|\|d_{k}\|\leq M_{3}\|d_{k}\|,$
where $M_{3}=L_{J}^{2}+M_{0}$. We conclude that
$-\nabla\phi(x_{k})^{T}d_{k}\geq\lambda_{k}\|d_{k}\|^{2}\geq\frac{\lambda_{k}}{M_{3}^{2}}\|\nabla\phi(x_{k})\|^{2}=\frac{1}{M_{3}^{2}}\|\nabla\phi(x_{k})\|^{3}$
if we take $\lambda_{k}=\|\nabla\phi(x_{k})\|$. Moreover,
$\|d_{k}\|=\|(J_{k}^{T}J_{k}+\lambda_{k}I)^{-1}\nabla\phi(x_{k})\|$ $\leq\|(J_{k}^{T}J_{k}+\lambda_{k}I)^{-1}\|\|\nabla\phi(x_{k})\|\leq 1,$
which shows that our “safeguard directions” are also gradient-related.
Now that it is proven that Algorithm 1 generates a sequence of directions $\{d_{k}\}$ that is gradient-related, using [4, Proposition 1.2.1], global convergence can be established.
Theorem 4.
Let $\{x_{k}\}$ be a sequence generated by Algorithm 1. Then, every limit point $\hat{x}$ of $\{x_{k}\}$ is such that $\nabla\phi(\hat{x})=0$.
Proof.
Let $K_{1}=\{k\in\mathbb{N}\mid\|J(x_{k}+d_{k})^{T}F(x_{k}+d_{k})\|\leq\vartheta\|J(x_{k})^{T}F(x_{k})\|\}$. If $K_{1}$ is infinite, it follows that $\|J(x_{k})^{T}F(x_{k})\|\rightarrow 0$, and therefore any limit point $\hat{x}$ of $\{x_{k}\}$ is such that $J(\hat{x})^{T}F(\hat{x})=0$, hence $\nabla\phi(\hat{x})=0$. Otherwise, if $K_{1}$ is finite, let us assume, without loss of generality, that $\|J(x_{k}+d_{k})^{T}F(x_{k}+d_{k})\|>\vartheta\|J(x_{k})^{T}F(x_{k})\|$, for each $k$, such that the step size is chosen to satisfy the Armijo condition. Since, by Proposition 1 and Remark 5, the directions of Algorithm 1 are gradient-related, it follows from [4, Proposition 1.2.1] that any limit point $\hat{x}$ of $\{x_{k}\}$ is a stationary point of $\phi(x)$. ∎∎
The next result establishes the connection between global and local convergence by showing that, under certain conditions, $\alpha_{k}=1$ for all $k$ sufficiently large and then Algorithm 1 becomes the “pure” LMMSS and the results of Section 4 are applicable.
Theorem 5.
Let $\{x_{k}\}$ be generated by Algorithm 1, using $x_{0}\in\mathbb{R}^{n}$ as the initial point. Suppose that Assumptions 1-3 and Assumption 4, with $r=1$, hold at the level set ${\cal L}_{0}=\{x\in\mathbb{R}^{n}\mid\phi(x)\leq\phi(x_{0})\}$. Moreover, assume that ${\cal L}_{0}$ is compact. Then, for every $k$ sufficiently large $\alpha_{k}=1$ and the sequence $\{\textup{dist}(x_{k},X^{*})\}$ converges to 0 quadratically.
Proof.
Note that by the definition of Algorithm 1 $\{\phi(x_{k})\}$ is decreasing. Starting from $x_{0}$, we conclude that the sequence $\{x_{k}\}$ remains in ${\cal L}_{0}$, a compact set. Therefore, there exists a convergent subsequence, as per Theorem 4, where every limit point of this subsequence is stationary. By removing this convergent subsequence, we can repeat this process successively for the remaining sequence, which remains in a compact set, thereby concluding that $\{x_{k}\}$ is the disjoint and enumerable union of convergent subsequences, each of which has a stationary point as limit.
Thus, for every $k$ sufficiently large, the norm of the gradient at $x_{k}$ will be sufficiently small. In other words, there is a positive integer $k_{0}$ such that
$\|J(x_{k})^{T}F(x_{k})\|\leq\frac{\omega^{3}\vartheta}{\tilde{C}L_{3}},\quad\text{for all}\quad k\geq k_{0}.$ (57)
For $k\geq k_{0}$, let $\bar{x}_{k+1}\in X^{*}$ such that $\|x_{k+1}-\bar{x}_{k+1}\|=\textup{dist}(x_{k+1},X^{*})$. Then, from (7), Assumption 3,(44) and (57), it follows that
$\|J(x_{k+1})^{T}F(x_{k+1})\|$ $=\|J(x_{k+1})^{T}F(x_{k+1})-J(\bar{x}_{k+1})^{T}F(\bar{x}_{k+1})\|$
$\leq L_{3}\|x_{k+1}-\bar{x}_{k+1}\|$
$=L_{3}\textup{dist}(x_{k+1},X^{*})\leq\frac{\tilde{C}}{\omega}L_{3}\textup{dist}(x_{k},X^{*})^{2}$
$\leq\frac{\tilde{C}L_{3}\|J(x_{k})^{T}F(x_{k})\|}{\omega^{3}}\|J(x_{k})^{T}F(x_{k})\|$ (58)
$\leq\vartheta\|J(x_{k})^{T}F(x_{k})\|,$
proving that $\alpha_{k}=1$ for each $k\geq k_{0}$. To obtain the quadratic convergence of $\{\textup{dist}(x_{k},X^{*})\}_{k\geq k_{0}}$ to 0, one may apply Theorem 3 when we have a diminishing rank or from Theorem 2, when $rank(J(x))=rank(J(x^{*}))\geq 1$. ∎∎
Remark 6.
We observe that in the case of zero residue a previous result [13, Theorem 3.1] assumed that the limit point $x^{*}$ is such that $F(x^{*})=0$. Also, in the case of unconstrained optimization, in order to prove a result similar to Theorem 5, the required assumption is that the Hessian at $x^{*}$ is positive definite; however, such assumption would imply that $x^{*}$ is an isolated stationary point. As we want to address the case of non-zero residue and possibly non-isolated stationary points, we considered the boundness of the level set instead of these other two conditions.
Remark 7.
In the scenario of diminishing rank, if Assumption 4 is satisfied with $10<r<$ 1, it is necessary to adjust the parameter $\lambda_{k}$ of Algorithm 1 to $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r}$. Furthermore, when this hypothesis holds, the above results follow with minor changes in some constants, and we can show there exists $k_{0}$ such that $\alpha_{k}=1,\forall k\geq k_{0}$ and that $\textup{dist}(x_{k},X^{*})$ goes to zero superlinearly.
6 Illustrative examples
The purpose of this section is to illustrate our convergence results by presenting some examples of non-zero residue nonlinear least-squares problems. In this section we use the shorthand notation (A1) – (A4) for Assumptions 14, respectively.
Example 1. Let us start with an example from [2, Example 5.1] with non-isolated minimizers (stationary-points). Consider the residual function
$F(x)=(x_{1}^{2}+x_{2}^{2}-1,x_{1}^{2}+x_{2}^{2}-9)^{T}.$
The set of minimizers (stationary-points) $X^{*}$ consists of points satisfying $x_{1}^{2}+x_{2}^{2}-5=0$, thus $\text{dist}(x,X^{*})=|\|x\|-\sqrt{5}|$. Moreover, for any $x^{*}\in X^{*}$, $\phi(x^{*})=\frac{1}{2}\|F(x^{*})\|^{2}=\frac{1}{2}(4^{2}+(-4)^{2})=16$.
It is straightforward to observe that
$J(x)=\left[\begin{array}[]{cc}2x_{1}&2x_{2}\\ 2x_{1}&2x_{2}\end{array}\right]$
has rank $1$ everywhere except for the origin. Moreover, as discussed in [2], Assumptions (A2), (A3) and (A4) with $r=1$ hold as well.
By considering LMMSS with the scaling matrix $L=[-1\quad 1]$, Assumption (A1) holds everywhere except for the line $x_{2}=-x_{1}$.
Hence, we are under conditions to apply the results of Section 4.1. As expected from Theorem 2, in Table 1, we can observe the local quadratic convergence of $\text{dist}(x_{k},X^{*})$ to zero starting from two distinct starting points in a neighborhood of $x^{*}=(0,\sqrt{5})^{T}\in X^{*}$.
Table 1: Stopping criterion $\|J_{k}^{T}F_{k}\|<10^{-8}$ with $x_{0}=(0,\sqrt{5}+0.03)^{T}$ as starting point (left), and $x_{0}=(0.01,\sqrt{5}-0.01)^{T}$ (right).
$k$ dist($x_{k},X^{*}$) $\left\lVert J_{k}^{T}F_{k}\right\rVert$ $k$ dist($x_{k},X^{*}$) $\left\lVert J_{k}^{T}F_{k}\right\rVert$
0 $1.3506\times 10^{-1}$ $1.2242$ 0 $4.4521\times 10^{-2}$ $3.9643\times 10^{-1}$
1 $1.7762\times 10^{-2}$ $1.5890\times 10^{-1}$ 1 $1.9821\times 10^{-4}$ $1.7729\times 10^{-3}$
2 $3.2402\times 10^{-7}$ $2.8982\times 10^{-6}$ 2 $3.8598\times 10^{-9}$ $3.4523\times 10^{-8}$
3 $1.1546\times 10^{-14}$ $1.0659\times 10^{-13}$
This example is also useful to illustrate the effect of using $L^{T}L\neq I$ and the globalization mechanisms of Algorithm 1.
Figure 1a depicts the trajectory of LMM ($L^{T}L=I$) and the one of LMMSS (with $L=[-1\quad 1]$) both starting from $x_{0}=(2,4)^{T}$. As we can see from the figure, while LMM dives perpendicularly to the set $X^{*}$ (observe the level curves in the figure), LMMSS trajectory is parallel to the null space of $L$, and they end up at different stationary points. As exposed in [5] for parameter identification problems, an advantage of using Algorithm 1 is that in a given problem with non-isolated solutions, we can choose $L$ that enforces a specific property. In this geometric example, by using $L=[-1,1]$, we aim for the method to prioritize stationary points where $\|L(x-x_{0})\|=|x_{2}-x_{1}-2|$ is as small as possible. In [5], the authors employed $L$ as discrete versions of derivative operators in order to promote smoothness in the obtained solutions for parameter identification problems in heat conduction. Moreover, Figure 1b enforces this aspect by showing that the presence of a well-thought choice of $L$ may conduce the iterations faster and in a more accurate manner than LMM, while it stimulates a desired property.
(a) Iterations and level curves of the objective function.
(b) Values of $\|J_{k}^{T}F_{k}\|$ until the stopping criterion is achieved.
Figure 1: Plot of iterations for LMM and LMMSS for Example 1, with starting point $(2,4)$.
The last aspect we shall discuss with this example is about the completeness condition and the global convergence mechanisms of Algorithm 1. As mentioned before, (A1) holds everywhere except for the line $x_{2}=-x_{1}$. LMMSS also tries to move parallel to the null space of $L$. Figure 2a illustrates what happens if we apply LMMSS without the safeguard Step 4, using $x_{0}=(-1,3)^{T}$. The displacements $x_{k}-x_{0}$ are in ${\cal N}(L)$, but the hyperplane $x_{0}+{\cal N}(L)$ does not intersect $X^{*}$. As a result, LMMSS minimizes $\phi(x)$ in $x_{0}+{\cal N}(L)$ approximating a point $\hat{x}$ where $\nabla\phi(\hat{x})\perp{\cal N}(L)$ so that the directions $d_{k}$ become orthogonal to $-\nabla\phi(x_{k})$ loosing the descent property and the algorithm crashes not satisfying the Armijo condition. Notice that $\hat{x}$ is such that $\hat{x}_{2}=-\hat{x}_{1}$ and (A1) fails. On the other hand, we use Algorithm 1 with the safeguard step, Step 4 corrects this behavior and when $x_{k}$ approaches $\hat{x}$ the classic LMM direction is employed instead; see Figure 2b. This allows the algorithm to move on and reach $X^{*}$.
(a) LMMSS without safeguard Step 4 in Algorithm 1.
(b) LMMSS with safeguard Step 4 in Algorithm 1.
Figure 2: Plot of iterations for LMM and LMMSS for Example 1, with starting point $(-1,3)$, and level curves of the objective function.
Example 2. Now we consider an example to illustrate the “diminishing rank” scenario. The residual function is given by
$F(x)=(x_{1}^{3}-x_{1}x_{2}+1,x_{1}^{3}+x_{1}x_{2}+1),$
which leads to $\phi(x)$ having an isolated global minimizer at $(-1,0)^{T}$ and a non-isolated set of local minimizers $\{x\in\mathbb{R}^{2}\mid x_{1}=0\}$ that we shall use as $X^{*}$ and, in this case, $\textup{dist}(x,X^{*})=|x_{1}|$.
This example was also analyzed in [2] where it was verified that assumptions (A2), (A3) and (A4) with $r=1$ hold in a neighborhood of $x^{*}=(0,2)^{T}$.
For assumption (A1), we point out that the null space of
$J(x)=\left[\begin{array}[]{cc}3x_{1}^{2}-x_{2}&-x_{1}\\ 3x_{1}^{2}+x_{2}&x_{1}\end{array}\right]$
is $\{0\}$ wherever $x_{1}\neq 0$ and it is $\text{span}\{(0,1)^{T}\}$ for $x\in X^{*}\setminus\{0\}$. Hence, it is clear that, regardless the choice of $L$, (A1) holds everywhere except for $X^{*}$.
In Table 2 and Figure 3 we report the behavior of LMM and LMMSS with $L=[-1\quad 1]$ starting in the neighborhood of $x^{*}=(0,2)^{T}$. From Table 2 the quadratic convergence rate is apparent for both methods. In Figure 3 we highlight the different limit points: while the classic LMM moves perpendicularly to $X^{*}$, LMMSS moves parallel to ${\cal N}(L)$, leading in this case to a stationary point with smaller $x_{2}$ coordinate.
Figure 3: Plot of iterations for LMM and LMMSS for Example 2, with starting point $(0.8,2.1)^{T}$, and level curves of the objective function. Limit points: $(1.4833\times 10^{-14},1.9915)^{T}$ for LMM and $(-7.5482\times 10^{-16},1.3377)^{T}$ for LMMSS.
Table 2: Values of dist($x_{k},X^{*}$) until the stopping criterion $\|J_{k}^{T}F_{k}\|<10^{-10}$ is achieved for Example 2 with $x_{0}=(0.8,2.1)^{T}$.
$k$ LMM LMMSS
1 $3.7143\times 10^{-1}$ $1.5307\times 10^{-1}$
2 $6.0270\times 10^{-2}$ $1.3438\times 10^{-2}$
3 $1.0055\times 10^{-3}$ $1.7991\times 10^{-4}$
4 $2.4684\times 10^{-7}$ $3.0097\times 10^{-8}$
5 $1.4833\times 10^{-14}$ $7.5482\times 10^{-16}$
Example 3. Let us consider $F(x)=(x_{1}^{2},x_{2}^{2},x_{1}+x_{2},1),$ the residual function. The corresponding least-squares function $\phi(x)$ has a unique global minimizer at $x^{*}=(0,0)^{T}$, which is the only stationary point of $\phi(x)$, $X^{*}=\{x^{*}\}$. The rank of
$J(x)=\left[\begin{array}[]{cc}2x_{1}&0\\ 0&2x_{2}\\ 1&1\\ 0&0\end{array}\right]$
is 2 for $x\neq x^{*}$ and it is 1 at $x^{*}$. Therefore the rank diminishes as we approach $X^{*}$ and (A1) holds everywhere, regardless $L$, except at $x^{*}$. Also, since $\|J(x)-J(y)\|\leq 2\|x-y\|$, (A2) is verified.
The distinguished feature of this example is that the error bound condition (A3) does not hold for every point in a neighborhood of $x^{*}$. In fact, by considering the sequence $x_{k}=(1/k,-1/k)^{T}$ it is not hard to show that
$\lim_{k\to\infty}\frac{\|\nabla\phi(x_{k})\|}{\textup{dist}(x_{k},X^{*})}=\lim_{k\to\infty}\frac{\|\nabla\phi(x_{k})\|}{\|x_{k}\|}=0,$
and thus (A3) cannot hold.
Nevertheless, for any $x\in\text{span}\{(1,1)^{T}\}={\cal N}(J(x^{*}))^{\perp}$, from the mean value theorem, we have
$\|\nabla\phi(x)-\nabla\phi(0)\|=\|\nabla^{2}\phi(\xi)x\|\geq 2\|x\|,$
where we used the fact that $(1,1)^{T}$ is an eigenvector of $\nabla^{2}\phi(\xi)$ and 2 is a lower bound for the corresponding eigenvalue. Therefore, in this example, starting at $x_{0}\in{\cal N}(J(x^{*}))^{\perp}$ and using $L$ such that ${\cal N}(L)={\cal N}(J(x^{*}))^{\perp}$, we force the sequence $\{x_{k}\}$ generated by LMMSS to stay in a subspace where (A3) holds. Also, with this choice of $L$, (A1) is satisfied everywhere, and since we have
$\|J(x)^{T}F(0)\|=0\leq K\|x-0\|^{2},$
(A4) with $r=1$ is fulfilled and quadratic convergence is achieved, as depicted in Table 3.
Table 3: Distances of $x_{k}$ to $x^{*}$ produced by LMMSS until the stopping criterion $\|J_{k}^{T}F_{k}\|<10^{-10}$ is achieved for Example 3 with $L=[-1\quad 1]$, for different starting points.
$k$ $x_{0}^{(1)}=(3,3)^{T}$ $x_{0}^{(2)}=(-2,-2)^{T}$
1 $2.0097$ $1.2571$
2 $8.0542\times 10^{-1}$ $3.8494\times 10^{-1}$
3 $1.5845\times 10^{-1}$ $2.4840\times 10^{-2}$
4 $1.9403\times 10^{-3}$ $7.6586\times 10^{-6}$
5 $3.6524\times 10^{-9}$ $2.6331\times 10^{-16}$
6 $7.4734\times 10^{-17}$
It is worth mentioning that even though (A3) does not hold for every $x$ in a neighborhood of $x^{*}$ in this example, a weaker Hölderian error bound condition [1] does hold. Precisely $\|\nabla\phi(x)\|\geq\|x\|^{3}$ or, in other words, $\nabla\phi$ is Hölder metrically subregular of order 1/3 (see [1, Definition 1]).
7 Final remarks
We have investigated convergence issues of Levenberg-Marquardt method with singular scaling (LMMSS) [5] for nonlinear least-squares problems with non-zero residue. Through a couple of modifications in the algorithm proposed in [5, Algorithm 1], mainly, the choice of Levenberg-Marquardt parameter and the line-search, we have established local and global convergence results under mild assumptions.
It was shown that LMMSS direction is well defined as long as the nullspace of the Jacobian and the nullspace of the singular scaling matrix $L^{T}L$ have trivial intersection (Assumption 1). Then, it was demonstrated local superlinear convergence rate under a local error bound condition on the gradient of the least-squares function (Assumption 3) and an assumption on the linearization error of the gradient (Assumption 4) which represents a measure of nonlinearity and residue size. These results were established for both the constant rank and diminishing rank cases.
By using an Armijo line-search and safeguard directions, it was proved that all limit points of the sequence generated by Algorithm 1 are stationary points of the least-squares function. We also presented some examples illustrating the assumptions and convergence results.
In future works we plan to consider Hölder metric subregularity [1] instead of the local error bound used here and address local convergence when a weaker form of Assumption 4 holds, namely when $\|J(x)^{T}F(z)\|={\cal O}(\|x-z\|)$ instead of $\|J(x)^{T}F(z)\|=o(\|x-z\|)$, following the lines of the analysis presented in [2].
Apart from tackling these questions, we aim to conduct computational experiments to test the algorithm, equipped with suitable stopping criteria, in certain inverse problems to assess its numerical behavior.
Acknowledgments
This work was supported by FAPESC (Fundação de Amparo à Pesquisa e Inovação do Estado de Santa Catarina) [grant number 2023TR000360]. The work of DG was partially supported by CNPq (Conselho Nacional de Desenvolvimento Científico e Tecnológico), Brazil [grant number 305213/2021-0].